What is the pressure inside a cylinder when the volume changes from 14.37 dm3 to 49.31 dm3 with an initial pressure of 9.021 x 102 kPa? The temperature of the cylinder remained constant at 22.06ºC.
1
Expert's answer
2018-12-07T01:49:19-0500

Answer on Question #83633, Chemistry/ General Chemistry

What is the pressure inside a cylinder when the volume changes from 14.37 dm3 to 49.31 dm3 with an initial pressure of 9.021 x 102 kPa? The temperature of the cylinder remained constant at 22.06ºC.

Solution

As temperature remained constant at 22.06C we can use Boyle's Law to find the final pressure incide a cylinder:

P1V1 = P2V2

where P1 = 9.021*102 kPa = 902.1 kPa

V1 = 14.37 dm3

V2 = 49.31 dm3

902.1*14.37 = P2*49.31

P2 = 262.9 kPa = 2.629*102 kPa

Answer: 2.629*102 kPa